OpenStudy (anonymous):

ΔRST is dilated about the origin, O, to create ΔR'S'T'. Point R is located at (3, 9), and point R' is located at (1.2, 3.6). Which scale factor was used to perform the dilation

OpenStudy (anonymous):

x-coordinates: R'/R = 1.2/3 = y-coordiantes: R'/R = 3.6/9 =